OpenStudy (anonymous):

Which of the following is the equation of a line parallel to 2y = 8x + 5 that passes through (2,7)?

OpenStudy (anonymous):

equation of any line parallel to it- \[8x-2y+k=0\] since the line passes through (2,7)- \[8*2-2*7+k=0\] \[k=-2\] so,the equation is- \[8x-2y-2=0\]